# Demystifying the Compression of Mixture-of-Experts Through a Unified Framework

### v2

@@ 4.1 Unified View for MoE Compression
Expert Trimming Block $\mathcal{T}\leftarrow\varnothing$
In the following sections, we will elaborate on how this framework incorporates existing methods and inspires us to identify new design space for compression.
@@ 4.2 Expert Slimming
Given that employing multiple experts in MoE significantly escalates parameters and inference costs, Expert Slimming, stemming from single-model compression techniques, targets the compression of individual expert weights $\boldsymbol{W}$ exclusively. We denote any efficient transformation function as $f(\cdot)$, which encompasses pruning $\boldsymbol{M}\odot\boldsymbol{W}$ and quantization $\mathrm{Quant}(\boldsymbol{W})$. Through the application of such functions, we reduce the redundancy within each expert and create several light-weighted slim experts, thus improving their intrinsic efficiency. However, it is important to note that Expert Slimming primarily focuses on compressing individual experts without addressing the redundancy across multiple experts. This brings us to another aspect of the unified framework, which deals with the structural redundancy present across the expert networks. To achieve comprehensive compression and efficiency, we need to consider both within-expert and across-expert redundancies, which will be explored further in the following sections.
4.3 Expert Trimming
Expert Trimming, which removes experts, is complementary to Expert Slimming. Therefore, we include Expert Trimming in our framework to provide a comprehensive view of MoE compression. The core operation of Expert Trimming involves updating the set of remaining experts denoted as $\mathcal{T}\leftarrow\mathcal{T}^{\prime}$, where $\mathcal{T}^{\prime}$ is a subset of the original expert indices $\mathcal{T}$. Specifically, Expert Drop updates the experts and their corresponding routing weights as follows: $\boldsymbol{E}\leftarrow\{\boldsymbol{E}_{i}\}_{i\in\mathcal{T}^{\prime}}$ and $\boldsymbol{G}\leftarrow\boldsymbol{G}_{i\in\mathcal{T}^{\prime}}$.
However, Expert Drop carries the risk of collapsing feature transformation. The absence of certain experts can lead to incorrect selections for given inputs, thereby degrading model performance [39]. Additionally, partially reducing experts can disrupt routing patterns, negatively impacting the model’s overall efficiency and effectiveness. Despite its benefits, Expert Drop still retains the costly computation within each expert and the complex communication between experts. These limitations highlight the need for further optimization of Expert Trimming within our unified framework. By systematically analyzing the redundancies and inefficiencies inherent in MoE models, our framework helps identify new design spaces for more efficient approaches, i.e., we propose extending beyond expert-level optimizations to identify new design spaces for efficiency improvements.
Figure 2: Illustration of Similarity Measurements in Layer Drop. Features for calculating $\boldsymbol{S}^{\text{(M)}}$ and $\boldsymbol{S}^{\text{(NM)}}$ are colored with red and blue, respectively.
We propose two novel techniques: Layer Drop and Block Drop. Layer Drop focuses on removing entire MoE layers, which significantly reduces both computation and communication overhead. Block Drop extends this concept by eliminating entire transformer blocks, including attention layers and MoE layers, within transformer models. These advanced techniques aim to streamline the model architecture, improve performance, and enhance overall efficiency.
Layer Drop.
Inspired by [53, 54], we consider a special scenario of Expert Drop where all experts are dropped ($\mathcal{T}\leftarrow\mathcal{T}^{\prime}=\varnothing$), effectively removing entire MoE layers. We refer to this approach as Layer Drop. To perform Layer Drop, we use a similarity-based metric where high similarity indicates high redundancy in transformation. One straightforward metric is the cosine similarity between the input $\boldsymbol{x}$ and the output $\boldsymbol{y}=\mathrm{MoE}(\boldsymbol{x})$:
$\boldsymbol{S}^{\text{(M)}}=\frac{\boldsymbol{x}\cdot\boldsymbol{y}}{||\boldsymbol{x}||_{2}\ ||\boldsymbol{y}||_{2}},~{}~{}\text{where}~{}~{}\boldsymbol{y}=\mathrm{MoE}(\boldsymbol{x}).$ (8)
However, this metric alone does not adequately capture the impact of the MoE layer within the context of a transformer block, which includes a layer normalization module ("Norm") [55] and residual connections [56]. To address this, we propose concurrently removing both the $\mathrm{MoE}$ and $\mathrm{Norm}$ layers. This approach ensures that the similarity metric more accurately reflects the combined functionality of these layers, allowing for a more precise identification of redundancy and a streamlined model architecture, as illustrated in Figure 1. By considering the similarity between the normalized input and output, we can better evaluate the necessity of the MoE layer in the overall architecture:
$\boldsymbol{S}^{\text{(NM)}}=\frac{\boldsymbol{x}^{\prime}\cdot\boldsymbol{y}^{\prime}}{||\boldsymbol{x}^{\prime}||_{2}\ ||\boldsymbol{y}^{\prime}||_{2}},~{}~{}\text{where}~{}~{}\boldsymbol{y}^{\prime}=\boldsymbol{x}^{\prime}+\text{MoE}(\text{Norm}(\boldsymbol{x}^{\prime})).$ (9)
Block Drop.
Within a transformer block, Layer Drop removes the MoE layers but retains the computation-costly attention layers [57, 58]. To address this issue, we further utilize the same similarity-based metrics to investigate whether the attention layer can be dropped without a significant performance drop. If feasible, this allows us to drop the entire block within MoE models, thus enhancing efficiency. We introduce Block Drop as an extension of Layer Drop, which also removes the attention layers.
Compared to Expert Drop, both Layer Drop and Block Drop focus on structures beyond expert-level. Additionally, we summarize the efficiency contributions of all the discussed Expert Slimming and Expert Trimming methods in Table 1, highlighting the unique advantages of each approach. Since Expert Trimming is complementary to Expert Slimming, they can be seamlessly integrated to compress both individual experts and high-level structures, further enhancing overall efficiency.
5 Experiments
In this section, we first evaluate Expert Slimming and Expert Trimming separately, then investigate their integration for further efficiency. Lastly, we discuss the potential of future works inspired by our proposed framework and experimental results. Implementation details are provided in Appendix A.
5.1 Expert Slimming

### v3

@@ 4.1 Overview
Expert Trimming Block Drop $\mathcal{T}\leftarrow\varnothing$
Pruning $\bm{M}\odot\bm{W}$
Expert Slimming Quantization Quant($\bm{W}$)
4.2 Expert Trimming
The core operation of Expert Trimming involves updating the set of remaining experts denoted as $\mathcal{T}\leftarrow\mathcal{T}^{\prime}$, where $\mathcal{T}^{\prime}$ is a subset of the original expert indices $\mathcal{T}$. Specifically, Expert Drop updates the experts and their corresponding routing weights as follows: $\bm{E}\leftarrow\{\bm{E}_{i}\}_{i\in\mathcal{T}^{\prime}}$ and $\bm{G}\leftarrow\bm{G}_{i\in\mathcal{T}^{\prime}}$.
However, Expert Drop carries the risk of collapsing feature transformation. The absence of certain experts can lead to incorrect selections for given inputs, thereby degrading model performance (Chen et al., 2022). Additionally, partially reducing experts can disrupt routing patterns, negatively impacting the model’s overall efficiency and effectiveness. Despite its benefits, Expert Drop still retains the costly computation within each expert and the complex communication between experts. These limitations highlight the need for further optimization of Expert Trimming to promote the efficiency. By systematically analyzing the redundancies and inefficiencies inherent in MoE models, we propose extending beyond expert-level optimizations to identify new design spaces for efficiency improvements.
Figure 2: Illustration of Similarity Measurements in Layer Drop. Features for calculating $\bm{S}^{\text{(M)}}$ and $\bm{S}^{\text{(NM)}}$ are colored with red and blue, respectively.
We propose two novel techniques: Layer Drop and Block Drop. Layer Drop focuses on removing entire MoE layers, which significantly reduces both computation and communication overhead. Block Drop extends this concept by eliminating entire blocks, including attention layers and MoE layers, within transformer models. These advanced techniques aim to streamline the model architecture, improve performance, and enhance overall efficiency.
Layer Drop
Inspired by Raposo et al. (2024); Elhoushi et al. (2024), we consider a special scenario of Expert Drop where all experts are dropped ($\mathcal{T}\leftarrow\mathcal{T}^{\prime}=\varnothing$), effectively removing entire MoE layers. We refer to this approach as Layer Drop. To perform Layer Drop, we use a similarity-based metric where high similarity indicates high redundancy in transformation. One straightforward metric is the cosine similarity between the input $\bm{x}$ and the output $\bm{y}=\mathrm{MoE}(\bm{x})$:
$\bm{S}^{\text{(M)}}=\frac{\bm{x}\cdot\bm{y}}{||\bm{x}||_{2}\ ||\bm{y}||_{2}},~{}~{}\text{where}~{}~{}\bm{y}=\mathrm{MoE}(\bm{x}).$ (8)
However, this metric alone does not adequately capture the impact of the MoE layer within the context of a transformer block, which includes a layer normalization module ("Norm") (Ba et al., 2016) and residual connections (He et al., 2015). To address this, we propose concurrently removing both the $\mathrm{MoE}$ and $\mathrm{Norm}$ layers. This approach ensures that the similarity metric more accurately reflects the combined functionality of these layers, allowing for a more precise identification of redundancy and a streamlined model architecture, as illustrated in Figure 2. By considering the similarity between the raw residual input and the aggregated output, we can better evaluate the necessity of the MoE layer in the overall architecture:
$\bm{S}^{\text{(NM)}}=\frac{\bm{x}^{\prime}\cdot\bm{y}^{\prime}}{||\bm{x}^{\prime}||_{2}\ ||\bm{y}^{\prime}||_{2}},~{}~{}\text{where}~{}~{}\bm{y}^{\prime}=\bm{x}^{\prime}+\text{MoE}(\text{Norm}(\bm{x}^{\prime})).$ (9)
Block Drop
Within a transformer block, Layer Drop removes the MoE layers but retains the computation-costly attention layers (Ribar et al., 2024; Zhang et al., 2023). To address this issue, we further utilize the same similarity-based metrics to investigate whether the attention layer can be dropped without a significant performance drop. If feasible, this allows us to drop the entire block within MoE models, thus enhancing efficiency. We introduce Block Drop as an extension of Layer Drop, which also removes the attention layers. Specifically, for the $i$-th block, we assess its importance score by evaluating the similarity between its inputs $\bm{x}_{l}$ and outputs $\bm{y}_{l}$. Compared to Expert Drop, both Layer Drop and Block Drop focus on structures beyond expert level, with the potential to further enhance the efficiency of MoE models.
@@ 4.3 Expert Slimming
Given that employing multiple experts in MoE significantly escalates parameters and inference costs, Expert Slimming, stemming from single-model compression techniques, targets the compression of individual expert weights $\bm{W}$ exclusively. We denote any efficient transformation function as $f(\cdot)$, which encompasses pruning $\bm{M}\odot\bm{W}$ and quantization $\mathrm{Quant}(\bm{W})$. Through the application of such functions, we reduce the redundancy within each expert and create several light-weighted slim experts, thus improving their intrinsic efficiency. However, it is important to note that Expert Slimming primarily focuses on compressing individual experts without addressing the redundancy across multiple experts. For maximum efficiency gains, Expert Slimming and Expert Trimming can be integrated to compress both individual experts and structured components. We summarize the efficiency contributions of all the discussed Expert Trimming and Expert Slimming methods in Table 1, highlighting the unique advantages of each approach.
5 Experiments on Expert Trimming
In this section, we evaluate the effectiveness of Expert Trimming techniques, starting with Expert Drop, and comparing it with our proposed methods, Layer Drop and Block Drop. Implementation details are provided in Appendix A.
Expert Drop: Performance Degradation with Limited Efficiency Gains
While experts are specific structures in MoE, not all experts hold equal significance. Figure 11 visualizes the distribution of expert-wise importance scores, highlighting this variability. To systematically drop experts at varying proportions, we conduct experiments using both layer-wise and global dropping approaches (see Appendix A.3). Given the importance of shared experts (Appendix E), we only dropped normal experts for DeepSeek-MoE-16B. Under both settings, Expert Drop causes consistent performance degradation. For example, dropping $25\%$ of experts in Mixtral-8$\times$7B results in a $23\%$ performance drop on the MMLU task. The efficiency improvement from Expert Drop is also marginal. For instance, dropping $12.5\%$ of experts results in less than a $1\%$ speedup, despite significant performance losses. More experimental results are available in Appendix F.
Figure 3: Evaluation of Expert Drop. We consider two strategies: layer-wise dropping (dotted lines) and global dropping (solid lines). “Random Guess” refers to randomly generating an output rather than using the model’s predictions, serving as a baseline to assess the extent of performance degradation.
Layer Drop: Comparable Performance with Greater Efficiency
To verify the feasibility of Layer Drop, we visualize feature similarity across different modules in Figure 4. This visualization shows a high level of similarity for features across the the MoE normalization module (Norm) and the MoE layer. In contrast, the low similarity for features across the MoE layer indicates the infeasibility of removing only MoE layers. Results from Figure 5 show that Layer Drop preserves performance within a wide range of compression ratio, e.g. $1\%$ performance drop on MMLU when dropping 8 layers for Mixtral-8$\times$7B, revealing significant redundancy in the MoE layers.
Block Drop: Further Optimizing Efficiency by Pruning Entire Transformer Blocks
[…]
Mistral-7B (Dense)
Method  ARC-C HellaSwag  MMLU  OBQA Average
Baseline 61.5 83.7 62.5 43.8 62.9
+ L4/32 53.2 77.7 61.7 40.0 58.2 (-4.7)
+ L8/32 36.7 33.6 53.3 30.6 38.6 (-24.3)
[…]
+ B8/32 40.0 63.9 60.0 30.6 48.6 (-14.3)
Mixtral-8$\times$7B (MoE)
Method  ARC-C HellaSwag  MMLU  OBQA Average
Baseline 59.4 84.0 67.9 46.8 64.6
+ L4/32 56.2 81.3 67.6 44.6   62.4 (-2.2)
+ L8/32 47.7 75.2 67.3 40.0   57.6 (-7.0)
[…]
+ B8/32 40.8 55.8 66.3 37.2 50.0 (-14.6)
While Layer Drop maintains the performance of the original models, it still preserves the computation-costly attention layers. To address this, Block Drop extends Layer Drop by removing whole transformer blocks, including both MoE and attention layers, further reducing computational and memory costs. Figure 8 visualizes block-wise similarity, where both Mixtral-8$\times$7B and DeepSeek-MoE-16B demonstrate high similarity between specific blocks. Based on this observation, we conduct the empirical study by varying the number of dropped blocks.
Surprisingly, as shown in Figure 6, the Mixtral-8$\times$7B maintains over $90\%$ of the original performance even after removing 5 blocks (over 7 billion parameters). Similar observations are also found in DeepSeek-MoE-16B, where 4 blocks can be removed when maintaining $90\%$ performance. Since Block Drop removes computationally expensive attention layers, it outperforms Layer Drop by a large margin in terms of both memory and inference cost, as illustrated in Figure 8.
On the other hand, Block Drop prunes attention layers along with their corresponding KV-Cache Pope et al. (2022). For instance, an input sequence with a batch size of 128 and a sequence length of 2048 results in 32GB of KV-Cache, which can be reduced by 5GB using Block Drop. Overall, by targeting higher-level structures, Layer Drop and Block Drop achieve substantial efficiency improvements while maintaining acceptable performance levels.
MoE Layers are More Redundant than Dense Counterparts
Since Layer Drop and Block Drop can also be applied to dense models, we take Mistral-7B, the corresponding dense model of Mixtral-8$\times$7B for comparison. Both models have the same depth and differ only in the FFN implementation, so we remove the same number of layers or blocks from each. When dropping an equal number of blocks, both MoE and dense models exhibit performance degradation. However, the MoE model suffers less performance drop under the same compression setting. For example, when dropping 8 MoE layers, the Mistral-7B receives a performance drop of 24.3, while Mixtral-8$\times$7B  only receives a drop of 7.0. This interesting finding highlights the higher redundancy in MoE layers, and further validates the effectiveness of applying Layer Drop and Block Drop to MoE models.
6 Visualization Examples of Layer Drop and Block Drop
In this section, we visualize the layer-wise similarity and the corresponding dropping order of MoE layers and blocks to investigate the varying levels of redundancy across different depths.
Since our similarity-based metrics depend on the hidden states of each block, the choice of data may influence feature similarity across layers. To investigate this, we conducted ablation studies on Mixtral-8$\times$7B, examining both the number of samples and the types of datasets used for feature extraction. This analysis helps us understand how data selection affects decisions regarding the dropping of layers or blocks. The results are presented in Figure 9.
(a) Similarities under different number of samples.
(b) Similarities under different datasets.
Figure 9: Influence of Data Choices on Feature Similarity. We measure the similarity among layers and blocks on Mixtral-8$\times$7B. (a) The similarity calculated using different number of samples from C4 Raffel et al. (2019). (b) The normalized similarity calculated using $1,024$ samples from different datasets, i.e., C4, Lima Zhou et al. (2023) and MetaMathQA Yu et al. (2023).
Robustness to Calibration Datasets
In Figure 9(a), we note that the feature similarity remains relatively stable across different layers as the sample size increases, indicating that Layer Drop and Block Drop maintain consistency regardless of the sample quantity. This confirms that using $128$ samples suffices for computing similarity, which is adopted for all our experiments. Similarly, Figure 9(b) shows that varying the datasets, from pretraining with C4 to instruction tuning with Lima and MetaMathQA, does not significantly alter the feature similarity. This demonstrates the resilience of Layer Drop and Block Drop to variations in data distribution.
(a) Layer Drop
(b) Block Drop
Redundant Deeper Layers
Figure 10 visualizes the remaining and dropped layers/blocks as the number of dropped modules increases. Both MoE architectures exhibit similar patterns in Layer Drop and Block Drop: initially, both models tend to drop the deeper layers, followed by the shallower ones. These findings are consistent with Xu et al. Men et al. (2024), which suggests that deeper layers tend to be more redundant.
[…]
Mixtral-8$\times$7B
Baseline 54.4T 87.7GB 59.4 84.2 84.0 67.9 46.8 83.8 70.4 75.6 71.5
[…]
+ E2/8 1.06×1.06$\times$1.06 × 54.4T 66.7GB 53.2 77.7 80.5 52.2 46.2 81.7 55.6 76.8 65.5
[…]
DeepSeek-MoE-16B
Baseline 11.7T 30.8GB 48.1 72.4 77.3 37.9 44.0 80.4 63.9 70.3 61.8
[…]
+ E16/64 1.06×1.06$\times$1.06 × 11.7T 23.9GB 45.0 67.1 75.6 31.8 42.2 80.2 59.9 70.0 59.0
[…]
7 Integration of Expert Trimming and Expert Slimming
Beyond Expert Trimming, another avenue for MoE compression is Expert Slimming, which targets individual expert compression. Techniques such as quantization and network pruning are among the most commonly employed methods. We provide a detailed comparison of network pruning and quantization in Appendix B, where the quantization outperforms both in performance and efficiency.
Since Expert Trimming and Expert Slimming focus on different aspects of compression, we further explore their potential integration. Given the superior average performance and practical efficiency of quantization, we use it for Expert Slimming. For Expert Trimming, we include all three methods to provide a complete comparison. The discussion on the order of applying these two compression techniques is provided in the Appendix D.
Quantization Preserves the Performance of Expert Trimming
As shown in Table 3, the integration of Expert Slimming and Expert Trimming significantly enhances overall efficiency. Quantization can be seamlessly combined with three different levels of dropping, achieving comparable performance. For instance, after quantization, the average performance of Layer Drop and Block Drop is nearly the same, maintaining more than $90\%$ of the performance of the original models.
The Integration Significantly Enhances Efficiency
In Table 3, the integration of Expert Trimming and quantization promotes efficiency by a large margin. Different Expert Trimming strategies showcase different advantages. Specifically, Expert Drop contributes to reducing memory usage, but its speedup is marginal. Layer Drop and Block Drop excel in speedup as illustrated in Figure 8, with Block Drop demonstrating both higher performance and greater speedup. Taking into account all settings, the combination of Block Drop and quantization offers the best efficiency with comparable performance: a 6.05×6.05$\times$6.05 × speedup with only $20.0$GB memory usage, while maintaining over $92\%$ of the performance on Mixtral-8$\times$7B, making it available to be deployed on a NVIDIA RTX 3090 GPU.
8 Post-Finetuning Recovers the Performance
While the discussed compression techniques maintains most of the performance of the original models, we further conduct post-finetuning to recover the degraded performance. Specifically, for comparison, we full-finetune DeepSeek-MoE-16B and corresponding compressed models on the Alpaca-GPT4 dataset (Peng et al., 2023) for 3 epochs using a learning rate of 8e-6 with 0.03 warmup ratio and cosine scheduling, where the global batch size is set to 32. As shown in Figure 4, the post-finetuning process significantly reduces the performance gap between the compressed models and the original models, e.g. narrowing it from 5.5% to 0.6% for the model following Block Drop.
Table 4: Performance of the DeepSeek-MoE-16B models finetuned after Expert Trimming. We mark the relative average performance loss of the compressed models compared to baselines in brackets.
DeepSeek-MoE-16B
Method SpeedUp FLOPs Memory ARC-C BoolQ HellaSwag MMLU OBQA PIQA RTE WinoGrande Average
Baseline 11.7T 30.8GB 48.1 72.4 77.3 37.9 44.0 80.4 63.9 70.3 61.8
+SFT 44.6 75.3 79.0 40.3 44.6 80.3 70.4 71.7 63.3
+ E16/64 1.06×1.06$\times$1.06 × 11.7T 23.9GB 45.0 67.1 75.6 31.8 42.2 80.2 59.9 70.0 59.0 (-2.8)
+SFT 44.4 74.0 78.6 38.5 45.8 79.6 65.7 70.1 62.1 (-1.2)
+ L4/28 1.14×1.14$\times$1.14 × 10.6T 26.6GB 39.5 70.2 67.6 35.2 40.4 75.8 48.4 65.7 55.3 (-6.5)
+SFT 42.1 78.9 75.2 40.8 43.4 77.6 71.1 69.5 62.3 (-1.0)
+ B4/28 1.16×1.16$\times$1.16 × 10.1T 26.4GB 40.3 71.3 69.0 36.2 37.8 75.8 51.6 68.0 56.3 (-5.5)
+SFT 43.2 78.2 75.0 40.4 43.8 76.8 74.0 70.2 62.7 (-0.6)
9 Conclusion
In this paper, we conducted a holistic study of MoE compression techniques, facilitating a systematic understanding of the efficiency issue of MoE and identifying the new design space to improve the performance further. Based on this study, we propose a comprehensive recipe that integrates Expert Slimming and Expert Trimming to further enhance efficiency. Our proposed methods and insights not only address current challenges but also set the stage for future advancements in the field of MoE.
Appendix A Implementation Details
A.1 Models and Datasets
Models
For our experiments, we employed Mixtral-8$\times$7B (Jiang et al., 2024) and DeepSeek-MoE-16B (Dai et al., 2024). Mixtral-8$\times$7B utilizes 8 experts for MoE layers and activates the top two for each input token. In contrast, DeepSeek-MoE-16B employs a dense FFN in the first block and utilizes two shared experts with an additional 64 experts within MoE layers in other blocks.
Datasets
For compression experiments, we used the C4 dataset (Raffel et al., 2019), with 128 samples and an input sequence length of 2,048, following the setup in (Sun et al., 2023; Lu et al., 2024; Lin et al., 2024; Frantar et al., 2022). To evaluate model performance, we report normalized zero-shot accuracy on the LM-harness benchmark, which includes multiple tasks: ARC-C (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2019), RTE (Wang et al., 2019), and WinoGrande (ai2, 2019). The evaluation code is based on EleutherAI LM Harness (Gao et al., 2023).
A.2 Implementation Details of Expert Slimming
Both Expert Slimming methods (i.e., pruning and quantization) require calibration data to estimate input statistics. To control this variable, we use $128$ samples from the C4 dataset (Raffel et al., 2019) as the calibration dataset for pruning. For quantization, we follow the default settings of GPTQ  [^1] https://github.com/AutoGPTQ/AutoGPTQ and AWQ  [^2] https://github.com/casper-hansen/AutoAWQ, using $128$ random samples from Alpaca (Taori et al., 2023) and Pile (Gao et al., 2020), respectively. We use the default group size $128$ for Mixtral-8$\times$7B and $64$ for DeepSeek-MoE-16B.
A.3 Implementation Details of Expert Drop
The Expert Drop compresses MoE by preserving only important experts $\{\bm{E}_{i}\}_{i\in\mathcal{T}^{\prime}}$ while removing others, where $\mathcal{T}^{\prime}$ is determined by the importance scores $\{\bm{S}(\bm{E}_{i})\}_{i\in\mathcal{T}}$. Following Muzio et al. (2024), we measure the importance scores through the averaged routing scores of a batched data $\mathcal{X}$, i.e., $\{\bm{S}(\bm{E}_{i})\}=\frac{1}{|\mathcal{X}|}\sum_{\bm{x}\in\mathcal{X}}\bm{G}_{i}(\bm{x})$, and consider two dropping strategies for Expert Drop: layer-wise dropping and global dropping.
$\mathcal{T}^{\prime(l)}=\{\bm{E}_{t}^{(l)}\},\quad\text{where}\quad\bm{S}(\bm{E}_{t}^{(l)})\in\mathrm{TopK}(\{\bm{S}(\bm{E}_{i}^{(l)})\}_{i=1}^{n},n^{\prime}).$ (10)
$\mathcal{T}^{\prime(l)}=\{\bm{E}_{t}^{(l)}\},\quad\text{where}\quad\bm{S}(\bm{E}_{t}^{(l)})\in\mathrm{TopK}\Big{(}\bigcup_{j=1}^{m}\{\bm{S}(\bm{E}_{i}^{(j)})\}_{i=1}^{n},n^{\prime}L\Big{)}.$ (11)
Appendix B Expert Slimming
Pruning: Comparable Performance with Deployment Challenges.
In Table 5, we evaluate representative pruning algorithms (i.e., Wanda (Sun et al., 2023), SparseGPT (Frantar & Alistarh, 2023)) on Mixtral-8$\times$7B and DeepSeek-MoE-16B. Since DeepSeek-MoE-16B utilizes both shared experts and normal experts, we conduct an ablation study on whether to prune shared experts, as discussed in Appendix E. We find that unstructured pruning preserves more than $95\%$ of performance. However, it is not compatible with existing hardware. Conversely, the hardware-friendly semi-structured pruning (i.e., $4$:$8$ and $2$:$4$ patterns) undergoes a significant performance drop. Nevertheless, according to Lu et al. (2024), semi-structured sparsity is ineffective in speeding up MoE models.